You are given 6 to 1 odds against tossing three heads with three coins, meaning you win $6 if you succeed and you lose $1 if you fail. Find the expected value​ (to you) of the game. Would you expect to win or lose money in 1​ game? In 100​ games? Explain.

Find the expected value​ (to you) for the game.

___ $___ (Type an integer or a decimal rounded to the nearest hundredth as​ needed.)

Would you expect to win or lose money in 1​ game?

A.

I would expect neither because the outcome of one game cannot be predicted.

B.

I would expect to​ win, because the probability of winning a game is greater than the probability of losing.

C

I would expect to​ win, because the expected value is positive.

D.I would expect to​ lose, because the expected value is negative.

Would you expect to win or lose money in 100​ games?

A. I would expect to​ win, because the expected value is positive.

B.I would expect to​ lose, because the probability of losing a game is greater than the probability of winning.

C. I would expect to​ win, because the probability of winning a game is greater than the probability of losing.

D. I would expect to​ lose, because the expected value is negative
Answer #1

Probability getting head in 3 tops es = 1/2 x 1 x 1 / - - - Probability for not getting head = 1-2/2 = 3 - 2 Expected value f